Which requires more work: lifting a 10 π‘˜π‘” sack of rice to a vertical distance of 2.0 π‘š or lifting a 5 π‘˜π‘” sack of rice to a vertical distance of 4.0 m


1
Expert's answer
2021-10-26T17:18:31-0400

Let's find the work done in both cases:

a) lifting a 10 π‘˜π‘” sack of rice to a vertical distance of 2.0 π‘š:


"W=Fh=mgh=10\ kg\times9.8\ \dfrac{m}{s^2}\times2\ m=196\ J."

b) lifting a 5 π‘˜π‘” sack of rice to a vertical distance of 4.0 m:


"W=Fh=mgh=5\ kg\times9.8\ \dfrac{m}{s^2}\times4\ m=196\ J."


As we can see from the calculations, both cases require the same amount of work.